The time intervals between successive barges passing a certain point on a busy waterway have an exponential distribution with mean 8 minutes. a) Find the probability that the time interval between two successive barges is less than 5 minutes. b) Find a time interval 𝑡 such that we can be 95% sure that the time interval between two successive barges will be greater than 𝑡.


1
Expert's answer
2022-01-17T18:51:43-0500

We have "\mu = 8," so "\lambda= 0.125."

(a) The probability is


"P(T < 5) =\displaystyle\int_{0}^{5}0.125e^{-0.125t}dt=-[e^{-0.125t}]\begin{matrix}
 5 \\
 0
\end{matrix}""=-e^{-0.125(5)}+e^{-0.125(0)}=0.4647"

(b) We require


"P(T >t) =\displaystyle\int_{t}^{\infin}0.125e^{-0.125x}dx"

"=\lim\limits_{A\to \infin}[-e^{-0.125x}]\begin{matrix}
 A \\
 t
\end{matrix}=e^{-0.125t}=0.95"

"t=-8\ln0.95\ min"

"t\approx 0.4103\ min=24.62\ s"